OpenStudy (anonymous):

Solve by elimination 5x-14y=22 -6x+7y=3

OpenStudy (jhannybean):

You have two equations, 5x - 14y = 22 -6x +7y = 3 Let's start by eliminating the y first. Multiply the second equation by 2, this will give you 5x - 14y = 22 -12x +14y = 6 _________________ Add up both these equations and solve for x first.

OpenStudy (daniel14):

still need help?